# TeV afterglow emission from a multi-component GRB jet using the kinetic approach

### v1

@@ Abstract
Recent years have seen a growing sample of TeV emission detections in gamma-ray burst afterglows, as well as an increasing role for structured jets in afterglow modelling. Using a kinetic approach, we show that the structure of an afterglow jet impacts its TeV emission, with jets where the energy falls off more sharply with angle showing a decrease in Inverse Compton (IC) peak flux relative to synchrotron peak flux. We use a modified version of the code katu, to which we have added adiabatic expansion and a fully self-consistent treatment of IC cooling both for the electron and photon populations. We compare our results to the semi-analytical code afterglowpy, finding a good agreement with our model except at early times off axis where the effects of baryon loading are important. We compare electron cooling in the cases where there is no IC cooling, Thomson cooling and an inclusion of Klein-Nishina effects, finding that the spectra can only be distinguished if the Compton potential is significantly increased. The smooth and gradual transition of our KN cooling also leads to a disparity in the cooling between our results and semi-analytical solutions based on asymptotic limits. Finally, we combine best fit parameters from afterglowpy to reproduce the light curves of GRB 170817A in our model. For our choice of parameters, we find that GRB 170817 would not have been detected in the TeV domain if seen on-axis, even by the upcoming Cherenkov Telescope Array Observatory.
@@
pubyear: 2023pagerange: TeV afterglow emission from a multi-component GRB jet using the kinetic approachC
@@ 1 Introduction
Numerical models are well placed to incorporate IC scattering by directly integrating over the Klein-Nishina cross-section and particle populations, without reference to $Y$ (Jones, 1968; Blumenthal & Gould, 1970). Such modelling is readily applied to kinetic codes, which have been used both outside (e.g. see Mastichiadis & Kirk 1997; Vurm & Poutanen 2009), and within GRB research (eg. Derishev 2021; Rudolph et al. 2023). In particular, one-zone kinetic modelling, with KN effects, has been done in works such as Petropoulou & Mastichiadis (2009), Pennanen et al. (2014) and Fukushima et al. (2017), and fitting of GRB afterglows with TeV emission has been done by Derishev & Piran (2021) to GRB 190114C, and to GRB 201216C by Huang (2022).
@@
In this paper, we combine the single-zone modelling of general IC scattering, such as in Derishev & Piran (2021), and the structured jet modelling of Ren et al. (2024) by modifying a kinetic code to run multi-component structured jets, with the inclusion of fully self-consistent IC scattering and cooling for the electron and photon populations. We used the blazar kinetic code Katu (Jiménez-Fernández & van Eerten 2020; Jiménez Fernández 2022), due to its readily available library of particle processes and implicit solver. To this, we added support for KN-corrected IC cooling and adiabatic expansion, and then built a shell model into the code to represent the fluid evolution. By running multiple concurrent simulations based off the energy distribution of the jet, we could then obtain a representation of a structured jet, and calculate the total observed flux with respect to the equal arrival time surface (EATS). Katu also has support for hadronic species and particle processes, though we limit this work to looking at electron and photon populations only (an example of leptohadronic modelling of GRB TeV emission can be found in Klinger et al. 2024).
@@ 2.1.1 Updates to Katu
$\frac{\partial n_{e}}{\partial t}\bigg{|}_{\rm adb}=\frac{n_{e}}{\tau_{\rm ad}}-\frac{\left(\hat{\gamma}-1\right)}{\tau_{\rm ad}}\frac{\partial}{\partial\gamma_{e}}(n_{e}\gamma_{e}),$ (2)
@@
$\frac{1}{\tau_{\rm ad}}=\dot{\gamma}-3\gamma\frac{{\dot{R_{\rm sh}\mkern-13.0mu}\mkern 13.0mu}{}}{R_{\rm sh}}.$ (3)
@@ 2.5 Observed Flux
We note in passing that the observed flux will not be sensitive to our (approximate) choice of $\tau_{\rm{esc}}$, given that $n_{\rm ph}\propto 1/\tau_{\rm esc}$.
@@ 4.2 Impact of different IC cooling regimes
We have plotted the combined synchrotron and IC cooling from Katu (solid) and from their model (dotted) in Fig. 10(b). Synchrotron-only cooling is also included from Katu (dashed), though both models have identical synchrotron cooling. Initially, the models diverge from the beginning by a few orders of magnitudes, with Katu converging more rapidly to the synchrotron dominant regime. As time progresses, we see both models start to agree at low energies, though Katu continues to show a less energetic and more gradual transition from the IC regime to the synchrotron regime, the latest lab time having the longest transition. In contrast, the model provided by McCarthy & Laskar shows asymptotic relations which sharply turnover at higher energies. Overall, while the models agree generally with the evolution of the IC to synchrotron transition, they differ on the position and graduation of the change.
@@ 6.2 Modelling IC scattering
Subsequently, when we look at the impact on the synchrotron regime in Fig. 9, our KN results shows a more gradual and smooth transition from the Thomson to KN regime, as compared to the cooling term used by McCarthy & Laskar (2024). We observe that the break-away from IC and synchrotron cooling to synchrotron-only cooling occurs at a lower frequency compared to what was found by McCarthy & Laskar (2024), which will be caused by the difference in electron cooling between the two models, as was shown in Fig. 10(b). The additional numerical treatment causes the cooling to consistently transition from IC to synchrotron dominated cooling earlier and smoother in energy space. Once this occurs, the KN spectra will converge with the synchrotron result, as in McCarthy & Laskar (2024), though in practise the IC scattered photons prevent this from happening in the actual observed spectrum.
@@ 7 Conclusions
  • IC cooling regimes – By comparing our numerical KN-corrected solution for the cooling of electrons to electron Thomson cooling and synchrotron-only spectra (we treat photons using the full KN cross-section throughout), we find noticeable differences in the synchrotron spectrum if the Compton potential, and by extension the Compton Y-parameter, is high, though between the KN-corrected spectrum and synchrotron-only spectrum the only key difference is in the spectral slope, not the flux level. We also find a smoother transition between the Thomson and KN regimes than in the analytical model presented by McCarthy & Laskar (2024), as the numerical approach is not reliant on asymptotic solutions of the Compton Y-parameter, but instead calculates the cooling from the smoothly changing KN cross-section.
@@ Appendix A Derivation of scaled injection term and adiabatic expansion
$\frac{\partial n_{e}}{\partial t}$ $=\frac{2-\hat{\gamma}}{\tau_{\rm ad}}n_{e}-\frac{\hat{\gamma}-1}{\tau_{\rm ad}}\gamma_{e}\frac{\partial n_{e}}{\partial\gamma_{e}}.$ (46)
where $\tau_{\rm ad}=n/\dot{n}$ is the adiabatic timescale in the fluid frame. This can be reorganized in the form of Eq. 2:
$\frac{\partial n_{e}}{\partial t}$ $=\frac{n_{e}}{\tau_{\rm ad}}-\frac{\hat{\gamma}-1}{\tau_{\rm ad}}\frac{\partial}{\partial\gamma_{e}}\left(n_{e}\gamma_{e}\right).$ (47)
@@
$\frac{1}{\tau_{\rm ad}^{{}^{\prime}}}$ $=\frac{\Delta n_{\rm ad}}{n\Delta t}=\frac{\dot{\gamma}}{\gamma}-3\frac{{\dot{R_{\rm sh}\mkern-11.0mu}\mkern 11.0mu}{}}{R_{\rm sh}},$ (56)
@@
$\frac{1}{\tau_{\rm ad}}=\dot{\gamma}-3\gamma\frac{{\dot{R_{\rm sh}\mkern-11.0mu}\mkern 11.0mu}{}}{R_{\rm sh}}.$ (57)
@@ Appendix B Updating the Kinetic solver
$\frac{\partial n_{e}}{\partial t}=Q_{e}+Q_{i}+\mathcal{L}+\frac{n_{e}}{\tau_{\rm ad}}-\frac{n_{e}}{\tau_{\rm esc}}-\frac{n_{e}}{\tau_{\rm dec}}\\ -\frac{\partial}{\partial\gamma_{e}}\left(\frac{d\gamma_{e}}{dt}n_{e}+({\hat{\gamma\mkern 2.0mu}\mkern-2.0mu}{}-1)\frac{1}{\tau_{\rm ad}}n_{e}\gamma_{e}\right).$ (59)
@@
$A$ $=\frac{2-{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}}{\tau_{ad}}-\frac{1}{\tau_{\rm esc}}-\frac{1}{\tau_{\rm dec}}-\frac{1}{\tau_{\rm acc}}-2S\gamma_{e}-\frac{\partial}{\partial\gamma_{e}}\left(\frac{d\gamma_{\rm ic}}{dt}\right),$ (63)
$B$ $=-\frac{\gamma_{e}}{\tau_{\rm acc}}-S\gamma_{e}^{2}-\frac{{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}-1}{\tau_{\rm ad}}\gamma_{e}-\frac{d\gamma_{\rm ic}}{dt}.$ (64)
We note that $\left(2-\hat{\gamma}\right)$ reappears for $n_{e}$ as $\gamma_{e}$ goes outside the partial derivative on the rightmost side.
@@
$A^{{}^{\prime}}=L+\frac{2-{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}}{\tau_{\rm ad}}$ $-\frac{1}{\tau_{\rm esc}}-\frac{1}{\tau_{\rm dec}}$ (67)
@@
$B^{{}^{\prime}}$ $=-\frac{1}{\tau_{\rm acc}}-\frac{{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}-1}{\tau_{\rm ad}}-S\gamma_{e}-\frac{1}{\gamma_{e}}\frac{d\gamma_{e}}{dt}.$ (69)

### v2

# TeV afterglow emission from a structured GRB jet using the kinetic approach

@@ Abstract
Recent years have seen a growing sample of TeV emission detections in gamma-ray burst afterglows, as well as an increasing role for structured jets in afterglow modelling. Using a kinetic approach, we show that the structure of an afterglow jet impacts its TeV emission, with jets where the energy falls off more sharply with angle showing a decrease in Inverse Compton (IC) peak flux relative to synchrotron peak flux. We use a modified version of the code katu, to which we have added adiabatic expansion and a fully self-consistent treatment of IC cooling both for the electron and photon populations. We compare our results to the semi-analytical code afterglowpy, finding a good agreement with our model except at early times off axis where the effects of baryon loading are important. We compare electron cooling in the cases where there is no IC cooling, Thomson cooling and an inclusion of Klein-Nishina effects, finding that the spectra can only be distinguished if the Compton potential is significantly increased. We obtain a similar cooling rate compared to semi-analytical solutions, with some small difference at the transition from IC to synchrotron dominated cooling. Finally, we use best-fit parameters determined by afterglowpy and re-scaled for our model to reproduce the light curves of GRB 170817A. For our choice of parameters, we find that GRB 170817 would not have been detected in the TeV domain if seen on-axis, even by the upcoming Cherenkov Telescope Array Observatory.
@@
pubyear: 2023pagerange: TeV afterglow emission from a structured GRB jet using the kinetic approachC
@@ 1 Introduction
Numerical models are well placed to incorporate IC scattering by directly integrating over the Klein-Nishina cross-section and particle populations, without reference to $Y$ (Jones, 1968; Blumenthal & Gould, 1970). Such modelling is readily applied to kinetic codes, which have been used both outside (e.g. see Mastichiadis & Kirk 1997; Vurm & Poutanen 2009), and within GRB research (eg. Derishev 2021; Rudolph et al. 2023). In particular, one-zone kinetic modelling, with KN effects, has been done in works such as Petropoulou & Mastichiadis (2009), Pennanen et al. (2014) and Fukushima et al. (2017), and fitting of GRB afterglows with TeV emission has been done by Derishev & Piran (2021) to GRB 190114C, GRB 201216C by Huang (2022) and GRB 221009A by Banerjee et al. (2024).
@@
In this paper, we combine the single-zone modelling of general IC scattering, such as in Derishev & Piran (2021), and the structured jet modelling of Ren et al. (2024) by modifying a kinetic code to run multi-component structured jets, with the inclusion of fully self-consistent IC scattering and cooling for the electron and photon populations. We used the blazar kinetic code Katu (Jiménez-Fernández & van Eerten 2020; Jiménez Fernández 2022), due to its readily available library of particle processes and implicit solver. To this, we added support for KN-corrected IC cooling and adiabatic expansion, and then built a shell model into the code to represent the fluid evolution. By running multiple concurrent simulations based off the energy distribution of the jet, we could then obtain a representation of a structured jet, and calculate the total observed flux with respect to the equal arrival time surface (EATS). Katu also has support for hadronic species and particle processes, though we limit this work to looking at electron and photon populations only (examples of leptohadronic modelling of GRB TeV emission can be found in Klinger et al. 2024; Banerjee et al. 2024).
@@ 2.1.1 Updates to Katu
$\frac{\partial n_{e}}{\partial t}\bigg{|}_{\rm adb}=-\frac{n_{e}}{\tau_{\rm ad}}-\frac{\left(1-\hat{\gamma}\right)}{\tau_{\rm ad}}\frac{\partial}{\partial\gamma_{e}}(n_{e}\gamma_{e}),$ (2)
@@
[…]
@@ 2.5 Observed Flux
We note in passing that the observed flux will not be sensitive to our (approximate) choice of $\tau_{\rm{esc}}$, given that $n_{\rm ph}\propto\tau_{\rm esc}$.
@@ 4.2 Impact of different IC cooling regimes
We have plotted the combined synchrotron and IC cooling from Katu (solid) and from their model (dotted) in Fig. 10(b). Synchrotron-only cooling is also included from Katu (dashed), though both models have identical synchrotron cooling. Initially, the models diverge from the beginning by around an order of magnitude, with Katu converging more rapidly to the synchrotron dominant regime. As time progresses, we see both models generally agree, with Katu having a smoother transition from the IC dominated to the synchrotron dominated regime, the latest lab time having the longest transition. In contrast, the model provided by McCarthy & Laskar shows asymptotic relations which show sharp breaks, but which generally agrees with Katu. Overall, we get good agreement between the two models, except at early times low energy, and with some small differences of cooling rates at the transitional phase between the IC and synchrotron cooling regimes.
@@ 6.2 Modelling IC scattering
Subsequently, when we look at the impact on the synchrotron regime in Fig. 9, our KN results shows a smoother transition from the Thomson to KN regime, as compared to the cooling term used by McCarthy & Laskar (2024). We observe that the Katu cooling term begins transitioning at lower energies compared to McCarthy & Laskar (2024), though shortly thereafter their cooling also turns over, and stays within an order of magnitude of the Katu result. The exception to this is the early time result, where we observe an order of magnitude difference before they converge in the synchrotron regime. Once the transition occurs, the KN spectra will converge with the synchrotron result, as in McCarthy & Laskar (2024), though in practise the IC scattered photons prevent this from being observed in the actual observed spectrum.
@@ 7 Conclusions
  • IC cooling regimes – By comparing our numerical KN-corrected solution for the cooling of electrons to electron Thomson cooling and synchrotron-only spectra (we treat photons using the full KN cross-section throughout), we find noticeable differences in the synchrotron spectrum if the Compton potential, and by extension the Compton Y-parameter, is high, though between the KN-corrected spectrum and synchrotron-only spectrum the only key difference is in the spectral slope, not the flux level. We also find a smoother, albeit similar, transition between the Thomson and KN regimes than in the analytical model presented by McCarthy & Laskar (2024), as the numerical approach is not reliant on asymptotic solutions of the Compton Y-parameter, but instead calculates the cooling from the smoothly changing KN cross-section.
@@ Appendix A Derivation of scaled injection term and adiabatic expansion
$\frac{\partial n_{e}}{\partial t}$ $=\frac{\hat{\gamma}-2}{\tau_{\rm ad}}n_{e}-\frac{1-\hat{\gamma}}{\tau_{\rm ad}}\gamma_{e}\frac{\partial n_{e}}{\partial\gamma_{e}}.$ (46)
where $\tau_{\rm ad}=-n/\dot{n}$ is the adiabatic timescale in the fluid frame. This can be reorganized in the form of Eq. 2:
$\frac{\partial n_{e}}{\partial t}$ $=-\frac{n_{e}}{\tau_{\rm ad}}-\frac{1-\hat{\gamma}}{\tau_{\rm ad}}\frac{\partial}{\partial\gamma_{e}}\left(n_{e}\gamma_{e}\right).$ (47)
@@
$\frac{1}{\tau_{\rm ad}^{{}^{\prime}}}$ $=\frac{\Delta n_{\rm ad}}{n\Delta t}=3\frac{{\dot{R_{\rm sh}\mkern-11.0mu}\mkern 11.0mu}{}}{R_{\rm sh}}-\frac{\dot{\gamma}}{\gamma},$ (56)
@@
[…]
@@ Appendix B Updating the Kinetic solver
$\frac{\partial n_{e}}{\partial t}=Q_{e}+Q_{i}+\mathcal{L}-\frac{n_{e}}{\tau_{\rm ad}}-\frac{n_{e}}{\tau_{\rm esc}}-\frac{n_{e}}{\tau_{\rm dec}}\\ -\frac{\partial}{\partial\gamma_{e}}\left(\frac{d\gamma_{e}}{dt}n_{e}+(1-{\hat{\gamma\mkern 2.0mu}\mkern-2.0mu}{})\frac{1}{\tau_{\rm ad}}n_{e}\gamma_{e}\right).$ (59)
@@
$A$ $=\frac{{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}-2}{\tau_{ad}}-\frac{1}{\tau_{\rm esc}}-\frac{1}{\tau_{\rm dec}}-\frac{1}{\tau_{\rm acc}}-2S\gamma_{e}-\frac{\partial}{\partial\gamma_{e}}\left(\frac{d\gamma_{\rm ic}}{dt}\right),$ (63)
$B$ $=-\frac{\gamma_{e}}{\tau_{\rm acc}}-S\gamma_{e}^{2}-\frac{1-{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}}{\tau_{\rm ad}}\gamma_{e}-\frac{d\gamma_{\rm ic}}{dt}.$ (64)
We note that $\left(\hat{\gamma}-2\right)$ reappears for $n_{e}$ as $\gamma_{e}$ goes outside the partial derivative on the rightmost side.
@@
$A^{{}^{\prime}}=L+\frac{{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}-2}{\tau_{\rm ad}}$ $-\frac{1}{\tau_{\rm esc}}-\frac{1}{\tau_{\rm dec}}$ (67)
@@
$B^{{}^{\prime}}$ $=-\frac{1}{\tau_{\rm acc}}-\frac{1-{\hat{\gamma\mkern 3.0mu}\mkern-3.0mu}{}}{\tau_{\rm ad}}-S\gamma_{e}-\frac{1}{\gamma_{e}}\frac{d\gamma_{e}}{dt}.$ (69)